The young modulus of elasticity is defined as E=F/A/e/a

Obtain the dimension of E where f= force A= area and l are lengths


1
Expert's answer
2021-07-05T08:41:16-0400
"E=\dfrac{FL_0}{A\Delta L}=\dfrac{M\dfrac{L}{T^2}L}{L^2L}=ML^{-1}T^{-2}."